Grade 12Algebra

Out of 1800 students in a school, 750 opted basketball, 800 opted cricket, and the remaining opted table tennis. If a student can opt only one game, find the ratio of
Number of students who opted for basketball to the total number of students.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Ratio of number of students who opted basketball to the total number of students
= 750 / 1800 = 25 / 60 = 5 / 12

Thanks